### v1

# Reservoir Computing as a Language Model

Abstract
Large Language Models (LLM) have dominated the science and media landscape duo to their impressive performance on processing large chunks of data and produce human-like levels of text. Nevertheless, their huge energy demand and slow processing still a bottleneck for further increasing quality while also making the models accessible to everyone. To solve this bottleneck, we will investigate how reservoir computing performs on natural text processing, which could enable fast and energy efficient hardware implementations. Studies investigating the use of reservoir computing as a language model remain sparse. In this paper, we compare three distinct approaches for character-level language modeling, two different reservoir computing approaches, where only an output layer is trainable, and the well-known transformer-based architectures, which fully learn an attention-based sequence representation. We explore the performance, computational cost and prediction accuracy for both paradigms by equally varying the number of trainable parameters for all models. Using a consistent pipeline for all three approaches, we demonstrate that transformers excel in prediction quality, whereas reservoir computers remain highly efficient reducing the training and inference speed. Furthermore, we investigate two types of reservoir computing: a traditional reservoir with a static linear readout, and an attention-enhanced reservoir that dynamically adapts its output weights via an attention mechanism. Our findings underline how these paradigms scale and offer guidelines to balance resource constraints with performance.
I Introduction
Modern sequence modeling tasks, such as language modeling and machine translation, have been dominated by attention-based architectures-most prominently the transformer family [1, 2, 3, 4, 5, 6]- which excel at capturing long-range dependencies by learning contextual representations through self-attention layers and feed-forward networks. However, transformer training and inference incur substantial computational and energy costs, often requiring specialized hardware and limiting accessibility with tighter resource budgets.
Reservoir computing (RC) offers an alternative. In RC, a large, fixed, recurrent “reservoir” projects inputs into a high-dimensional state space, and only a lightweight readout layer is trained. This design dramatically reduces training time and energy consumption and can be implemented efficiently in software or on analog hardware substrates. Recent work has extended classical RC with neural programming techniques to shape reservoir dynamics and introduced attention mechanisms at the readout stage to adaptively weight reservoir states [7]. RC has been shown to excel in time-series prediction [8, 9, 10], capitalizing on the mentioned large, fixed random reservoir for capturing nonlinear dynamics with minimal training overhead [11, 12, 13, 14]. The charm of reservoir computing lies in its versatile implementation via a wide range of physical substrates [15], including quantum systems as potential RC candidates [16, 17, 18].
Recently, more expressive control over reservoir behavior has been introduced via neural programming paradigms, further expanding their utility in symbolic computation and structured sequence modeling [19]. While prior work has explored the use of reservoir computing for text classification tasks [20] and even demonstrated hardware-oriented implementations for language learning [21], the use of reservoir computing explicitly as a language model—i.e., for generative next-token prediction—remains largely unexplored. Reservoir computing has an advantage of simplicity, since only a small readout layer must be trained, while transformers require end-to-end training of embeddings, multi-head self-attention layers, and feed-forward networks. However, transformers can achieve state-of-the-art performance in language tasks by scaling to millions (or billions) of parameters [4].
In this study, we present a unified framework, applying both approaches to character-level sequence prediction on a small corpus of Shakespeare text inspired by recent minimalist transformer implementations such as Karpathy’s NanoGPT [22], which demonstrate how compact models can still achieve strong performance on character-level language modeling tasks. We vary reservoir size in the classic RC for a linear readout and RC size and attention layer size for an attention-enhanced readout [7], while varying hidden dimension and number of layers in transformers, obtaining multiple configurations spanning from thousands up to hundred thousand of trainable parameters. By measuring cross-entropy loss across these diverse setups, we expose fundamental trade-offs in resource usage and generalization performance.
Input Textshak$\begin{pmatrix}0.2\\ \vdots\\ 0.8\end{pmatrix}$$\begin{pmatrix}0.3\\ \vdots\\ 0.9\end{pmatrix}$$\begin{pmatrix}0.5\\ \vdots\\ 0.6\end{pmatrix}$$\begin{pmatrix}0.7\\ \vdots\\ 0.1\end{pmatrix}$TimeReservoirAtt-Enhanced ReservoirTransformer$\begin{pmatrix}w_{1}\\ w_{2}\\ \vdots\\ w_{m}\end{pmatrix}$ Output Layer$\begin{pmatrix}p_{a}=0.1\\ p_{b}=0.01\\ p_{c}=0.12\\ \vdots\end{pmatrix}$ $H(y,\hat{y})=-\sum_{i}y_{i}\log\hat{y}_{i}$ Cross Entropy – PredictionVector Embedding
Figure 1: Diagram illustrating the reservoir computing model with vector embedding, reservoir processing (Att-Enhanced Reservoir, Transformer), output layer, and predicted probability vector for the next letter. The vector embedding is trained for the transformer case, while being randomly initialized for the reservoirs. The predicted probabilities (e.g., $p_{A}$ for letter a) are used to compute the cross entropy loss $H(y,\hat{y})=-\sum_{i}y_{i}\log\hat{y}_{i}$, which drives the error feedback.
II Task Formulation and Dataset
We tackle a character-level next-character prediction task. Let $\mathbf{X}=\{x_{1},\dots,x_{T}\}$ represent a fixed-length input sequence of characters, and let $y_{T+1}$ be the next character to predict. We treat the problem as a multi-class classification over the vocabulary $\mathcal{V}=59$. The dataset is a 9-million-character corpus of Shakespeare text inspired by recent minimalist transformer implementations such as Karpathy’s NanoGPT [22], which we split into $6$ shards: $5$ shards for training and $1$ shard for testing. Each shard is processed at the character level, forming sequences of length $32$ for the input, plus the next character as the target. Before training the text corpus is preprocessed by lowercasing all letters to achieve a smaller vocabulary $\mathcal{V}=59$. Training optimizes a standard cross-entropy loss over the predicted logits versus the correct next character. We track training and test losses per shard and epoch.
III Background and Model Architectures
III.1 Traditional Reservoir Computing
The echo state implementation of a classic reservoir computer consists of three main components: an input-to-reservoir weight matrix $\mathbf{W}_{\mathrm{in}}\in\mathbb{R}^{N\times d}$, a fixed recurrent reservoir matrix $\mathbf{W}_{\mathrm{res}}\in\mathbb{R}^{N\times N}$, and a readout layer $\mathbf{W}_{\mathrm{out}}\in\mathbb{R}^{\mathcal{V}\times N}$ [11]. Here $d$ is the dimension of the chosen input-embedding vector for the input letter, $N$ the number of nodes in the reservoir, and $\mathcal{V}$ the size of the vocabulary. The reservoir evolves its hidden states $\mathbf{r}_{t}\in\mathbb{R}^{N}$ as
$\mathbf{r}_{t}=\tanh(\mathbf{r}_{t-1}\,\mathbf{W}_{\mathrm{res}}+\mathbf{x}_{t}\,\mathbf{W}_{\mathrm{in}}).$
Here, $\mathbf{x}_{t}\in\mathbb{R}^{d}$ is an embedding vector of the current input character. In the simplest (traditional) reservoir setting, only the readout layer $\mathbf{W}_{\mathrm{out}}$ is trained:
$\bar{\mathbf{y}}_{t}=\mathbf{r}_{t}\,\mathbf{W}_{\mathrm{out}}.$
where $\mathbf{r}_{t}$ is the reservoir state at time $t$. Typically, $\mathbf{W}_{\mathrm{out}}$ is learned using ridge regression, which minimizes the mean squared error with an added $\ell_{2}$ regularization term to prevent overfitting and improve generalization. However, in our experiments, we employ a cross-entropy loss function, necessitating gradient-based optimization methods for training $\mathbf{W}_{\mathrm{out}}$. Moreover, ridge regression can become computationally intensive as the number of training samples and reservoir size increase, due to the inversion of large matrices. By using gradient-based methods, we circumvent these scalability issues.
Table 1: Trainable Parameter Counts for the classic reservoir, AERC, and Transformer Architectures
Approach Configuration Parameters
Transformer Hidden Size $d_{h}=64$, Heads $h=4$, Layers $L=4$ $15067$
Hidden Size $d_{h}=72$, Heads $h=8$, Layers $L=8$ $30299$
Hidden Size $d_{h}=128$, Heads $h=8$, Layers $L=8$ $45083$
Hidden Size $d_{h}=356$, Heads $h=8$, Layers $L=8$ $105275$
Hidden Size $d_{h}=256$, Heads $h=16$, Layers $L=16$ $155803$
Reservoir Reservoir Size $N=250$ $14809$
Reservoir Size $N=500$ $29559$
Reservoir Size $N=750$ $44309$
Reservoir Size $N=1750$ $103309$
Reservoir Size $N=2600$ $153459$
Att Reservoir Reservoir Size $N=75$, Hidden Size $H=13$ $15464$
Reservoir Size $N=75$, Hidden Size $H=19$ $31124$
Reservoir Size $N=100$, Hidden Size $H=20$ $45259$
Reservoir Size $N=150$, Hidden Size $H=25$ $102809$
Reservoir Size $N=160$, Hidden Size $H=30$ $155459$
III.2 Attention-Enhanced Reservoir Computing
Incorporating an attention mechanism into a RC can significantly improve its performance by making the output weights adaptable to the reservoir states, although increasing the number of trainable weights relative to the reservoir size [7]. Instead of using fixed output weights as in traditional RC, the attention-enhanced reservoir computing (AERC) uses dynamically computed attention weights that vary with inputs.
For each data point $l$ fed into the reservoir, the corresponding attention weights $\mathbf{W}_{\text{att},l}\in\mathbb{R}^{H_{o}\times N}$ are calculated. These attention weights are derived using a small neural network $F$ with trainable parameters $\mathbf{W}_{\text{net}}$, based on the reservoir states $\mathbf{r}_{l}$:
$\mathbf{W}_{\text{att},l}=F(\mathbf{W}_{\text{net}},\mathbf{r}_{l}).$
The weights are used to project the reservoir state onto a small hidden state $\mathbf{r_{o}}_{l}\in\mathbb{R}^{H_{o}}$:
$\mathbf{r_{o}}_{l}=\mathbf{W}_{\text{att},l}\mathbf{r}_{l}.$
This intermediate output vector is then mapped onto the final logits of dimensions $\mathcal{V}$ via a final linear matrix $\mathbf{W}_{out}\in\mathbb{R}^{\mathcal{V}\times H_{o}}$:
$\bar{\mathbf{y}}_{l}=\mathbf{W}_{out}\mathbf{r_{o}}_{l}.$
This choice of architecture with an intermediate layer is done to reduce the number of parameters, resulting from the 3-d tensor that maps from the attention layer in the $F(\mathbf{W}_{\text{net}},\mathbf{r}_{l})$ to the attention weights $\mathbf{W}_{\text{net}}$ mapping from a vector to a matrix.
Our goal is to refine $\mathbf{W}_{\mathrm{net}}$ so that the AERC can effectively model complex temporal dependencies in the input data. In this work, the neural network component consists of a single hidden layer with a ReLU activation function. After the training phase, the attention mechanism computes new attention weights at each time step. These dynamically updated attention weights allow the system to adapt to the evolving input, improving performance in complex dynamical systems.
III.3 Transformer Architectures
Transformers [1] are fully trainable neural networks that have achieve state-of-the-art performance on many sequence modeling tasks. The input sequence of characters is mapped to embeddings of dimension $d$. A stack of $L$ layers follows, each containing a multi-head self-attention block and a feed-forward sub-block. Attention heads learn to focus on relevant parts of the input sequence at different positions. The feed-forward layers, typically parameterized by matrices of dimension on the order of $d\times\alpha d$ (where $\alpha$ is a constant expansion factor), refine the representation. The complexity of the transformer grows with $L$ (the number of layers) and $d$ (the embedding dimension). The output is passed through a final projection to logits of size $\mathcal{V}$, the vocabulary dimension. All parameters (including embedding, attention, feed-forward, and output projection) are learned end-to-end, giving transformers a large capacity to capture long-range dependencies. However, this capacity often comes at a significant computational and memory cost, particularly for large $L$, $d$ and long sequences.
IV Architectures
The number of trainable parameters depends strongly on the chosen architecture and hyperparameters. We focused on choosing hyperparameters, such that the number of trainable parameters is comparable between all 3 different models. Table 1 summarizes the number of trainable parameters for each architecture across different configurations. With a vocabulary size of $V=59$ characters due to the Shakespeare dataset used in our experiments we arrive at architectures that range from around 15000 trainable parameters to around 150000.
For all three models, we take an embedding dimension of $d=16$. For the Transformer model we vary the hidden size $d_{h}$, the number of heads $h$, and the number of layers $L$. For the classic reservoir and AERC configurations, we vary the reservoir size $N$, the hidden dimensions of the attention layer $H$, and set the final layer $H_{o}=H$ to be equal to the attention layers hidden dimension.
Input Text to be, or not clear born have the poor my turn with my soul. romeo what when queen let this shall let the sbatter did the death, well the part, that your grief is turned, we hard me, like on an i when a true, and her-promain the commone to with you this lodged prithee to be, or not as a find. hele to where i wheress thieve hot may come, and you the sease, thy agains for your gon. pring thy worney that light and he winder, eyes was from three like io be good reator, but have them beenes of be strengler. i have such a foglingers widdly, good time i, btund. to be, or not my could now richment. lodovico-ling, as yield make one in their if younger, farewell. duchall, then, my great, teater these is and that beauty brow man, for here is the lord of play birthing on the had your and, but i’ll gench i heart that this mans colors well TransformerReservoirAERCVector Embedding
Figure 2: Comparison of language models: Transformer, Reservoir, and AERC with around 155k parameters (biggest model in Table 1) run in auto-regressive mode with the initial seed ”to be, or not” showing the generated text.
V Training
We train the parameters of all three architectures to predict the next character in a sequence by minimizing the cross-entropy loss:
$\mathbf{W}_{\mathrm{net}}^{(s+1)}=\mathbf{W}_{\mathrm{net}}^{(s)}-\gamma\nabla F\big{(}\mathbf{W}_{\mathrm{net}}^{(s)},\mathbf{R}\text{/}\mathbf{X}\big{)},$ (1)
where $s$ indexes the training batches, $\gamma$ is the learning rate, and in the two reservoir computer cases $\mathbf{R}$ denotes the reservoir states and in the Transformer case $\mathbf{X}$ the input sentence embedding. One batch corresponds to a subset of the training data used in each update step. The cross-entropy loss is defined as
$H(y,\hat{y})=-\sum_{i}^{\mathcal{V}}y_{i}\log\hat{y}_{i}$ (2)
where $\mathcal{V}$ is the number of output classes, $\hat{y}_{i}$ are the output logits, and $y_{i}$ is the true one-hot encoded label. We apply the softmax function to obtain output probabilities and compute the cross-entropy loss accordingly, a standard formulation in classification tasks [23]. All trainable weights are updated using backpropagation [24] and optimized using the Adam optimizer, a widely adopted variant of stochastic gradient descent [25].
VI Methodology
We divide the Shakespeare corpus into $6$ shards. Five shards serve as training data, and one shard is used for testing. We sample sequences of length $32$, i.e. 32 characters as input, and train on a batch size of $1024$ with Adam at a learning rate $1\times 10^{-4}$. The traditional reservoir, the attention-enhanced reservoir, and the transformer models are trained entirely by backpropagation in a consistent framework. We measure cross-entropy on the test shard as our main performance metric. We also assess closed-loop generation by feeding the predicted character back as input, measuring distributional metrics and repetitive patterns. Our parameter sweep includes varying the reservoir size or the neural network size in the attention reservoir, as well as varying the hidden dimension $d$ and number of layers and heads $L$ in the transformer. To save compute power, we first compute all the reservoir responses in one shard and keep them in memory, after which a few epochs on this shard are run before the next shard is processed. This procedure is repeated multiple times for all shards.
VII Results
We now present the performance of the trained models and analyze the training process. Figure 2 shows a representative example of text generated by each of the three models, all configured with approximately 155,000 trainable parameters—corresponding to the largest models listed in Table 1. While the generated text mimics the style of the Shakespearean corpus, the limited model capacity and small vocabulary size lead to inconsistencies and occasional errors. Nonetheless, the examples demonstrate that all three models successfully learn character-level prediction distributions that capture essential patterns in the dataset.
To quantitatively assess model performance, we report training and testing losses across epochs and data shards for all five model sizes described in Table 1. The results are visualized in Figure 3.
Figure 3: Training and testing loss over accumulated number of shared epochs for the Transformer, Reservoir, and Attention-Enhanced Reservoir Computer for the 5 different complexity models of Table 1. The thin light-grey dashed vertical lines indicate a shard change, while the black dashed lines show one complete epoch of all shards. Every shard was run for 5 shard epochs until the next one was processed.
Figure 3 presents training (top row) and testing (bottom row) losses for the five model sizes listed in Table 1, ordered from smallest (turquoise) to largest (grey). Model sizes range from approximately 15,000 to 155,000 trainable parameters.
Training is performed in shards: for each shard, reservoir responses are precomputed and stored in memory. Multiple training epochs are then performed on each shard before proceeding to the next. This approach mimics the inherent speed of reservoir computing, which theoretically operates with near-instantaneous dynamics. Vertical light-gray dashed lines mark shard transitions, while thicker dark-gray dashed lines denote full passes over the dataset (epochs). Due to this sharding structure, we observe periodic spikes in both training and testing losses—particularly in training—caused by model fine-tuning on individual shards. While increased stochasticity might reduce overfitting and improve convergence by avoiding poor local minima, our configuration of six shards and five epochs per shard balances computational efficiency and generalization. This setup tries to maximize the benefits of precomputing reservoir states while mitigating overfitting through moderate stochasticity.
Analyzing the loss curves for the Transformer, the classical Reservoir Computer (RC), and the Attention-Enhanced Reservoir Computer (AERC), we observe that loss variance across different model sizes is more pronounced in the RC and AERC than in the Transformer. As expected, the Transformer achieves the best overall performance, with a minimum test loss of 1.67 for the largest configuration. However, we note that the classical RC achieves a test loss as low as 2.01, and the AERC reaches an intermediate value of 1.73—indicating that both RC-based models are competitive despite their relative simplicity.
Finally, the Transformer exhibits the smallest training–testing loss gap, reflecting better generalization and reduced overfitting. In contrast, both reservoir-based models show a greater tendency to overfit, likely due to the fixed nature of their reservoirs and the limited capacity of their readout layers.
An important direction for further research is to investigate scaling laws for reservoir computing in the context of large language model (LLM) prediction. These results would be consistent with known scaling laws of transformer architectures, where performance improves predictability with increased model size and dataset scale [26]. Understanding how performance scales with reservoir size and model capacity could offer valuable insights into the potential of RC-based architectures in more complex sequence modeling tasks. As computational resources improve, we plan to explore the behavior of both classical and attention-enhanced reservoir computers at larger scales, assessing their viability as alternatives to conventional deep learning models.
VII.1 Evaluation Metrics and Analysis
Next we investigate the long-term text similarity performance of the three architectures for auto-regressive generation. To assess the performance of the different model architectures in generating longer text, we employ the N-gram overlap to quantify performance beyond the loss of cross entropy. This metric quantifies the aspect of similarity to the reference data and distributional alignment and is inspired by classical N-gram-based evaluation metrics such as BLEU [27], which assess the proportion of overlapping sequences between generated and reference text. The following subsection details the metric.
VII.1.1 N-gram Overlap
N-gram Overlap measures the similarity between the generated text and a reference text by computing the fraction of unique n-grams in the generated text that also appear in the reference text.
[…]
where $G_{n}$ is the set of unique n-grams in the generated text, and $R_{n}$ is the set of unique n-grams in the reference (test) text.
VII.1.2 Mathematical Formulation
Given a generated text $G=\{g_{1},g_{2},\dots,g_{M}\}$ and the reference text $R=\{r_{1},r_{2},\dots,r_{N}\}$ we extracted the n-grams as:
$G_{n}=\left\{(g_{i},g_{i+1},\dots,g_{i+n-1})\mid 1\leq i\leq M-n+1\right\},$
$R_{n}=\left\{(r_{j},r_{j+1},\dots,r_{j+n-1})\mid 1\leq j\leq N-n+1\right\}.$
The overlap-n metric is then given by:
[…]
This metric assesses how well the model captures the n-gram patterns present in the reference data. A higher overlap indicates that the generated text shares more common sequences with the reference, suggesting better generalization to the learned patterns.
VII.2 N-Gram Similarity
The results of the 7-gram and 8-gram performance are shown in Fig. 4. The evaluation metrics reveal that the classic reservoir approach tends to have the lowest overlap and thus the lowest generalization. Transformers and Attention-Enhanced Reservoirs demonstrate on par diversity and distributional alignment, making them well-suited for tasks like text generation. Their ability to maintain high performance underscores their superior performance, due to the inclusion of neural networks.
The traditional reservoir, though the simplest, can struggle with capturing complex dependencies and maintaining distributional alignment, as evidenced by its lower distinct-n scores. However, its simplicity remains valuable in scenarios where an easy hardware implementation is prioritized.
The attention-enhanced reservoir improves upon the traditional model by incorporating dynamic readout weights, thereby enhancing its ability to capture more complex patterns without incurring the full computational burden of transformers. This architecture serves as a middle ground, offering improved performance over the traditional reservoir, while offering reducing complexity in the neural network architecture.
VII.3 Inference Time Comparison
Figure 4: 7-gram and 8-gram overlap for the Transformer, Reservoir, and Attention-Enhanced Reservoir Computer over the number of trainable parameters.
Figure 5: Training and inference time for a longer text generation over the number of trainable parameters for the Transformer, Reservoir, and Attention-Enhanced Reservoir Computer. The numbers above the line are from a fit applied via $y=\alpha\log_{10}(x)$, where $x$ is the number of trainable parameters, $y$ is the training time and $\alpha$ the slope.
As a final performance metric, we evaluate the training and inference time costs for all three model architectures. Figure 5 displays the time (in seconds) required for both training and inference, plotted against the number of trainable parameters in each model. Inference time refers to generating a sequence over a fixed data chunk or producing longer text samples.
Notably, we exclude the computational cost of the reservoir dynamics themselves. We assume that the reservoir computation can be offloaded to physical substrates such as photonic circuits, which have been experimentally shown to operate at extremely high speeds [28]. Such hardware would operate at time scales several orders of magnitude faster than the subsequent postprocessing by the trained layers, justifying its omission in this comparison.
In Figure 5, solid lines indicate training time, while dashed lines represent inference time. The color scheme distinguishes between the three model types: classic reservoir (blue), attention-enhanced reservoir (orange), and transformer (green, red, and purple). It is evident that reservoir-based models exhibit substantially lower computational times, and their scalability with respect to parameter count is more favorable than that of transformers. The numbers above the line are from a fit applied via $y=\alpha\log_{10}(x)$, where $x$ is the number of trainable parameters, $y$ is the training time and $\alpha$ the slope. From the graphs itself and the slope $\alpha$ we can see that reservoir based training and inference time is by magnitudes of order faster. This performance advantage arises because the reservoir handles the computationally expensive sequence processing, whereas in transformers, attention mechanisms scale quadratically with sequence length. These results suggest that using a reservoir for data preprocessing can significantly accelerate language model performance without incurring the high time cost associated with full transformer architectures.
VIII Conclusion
We presented a unified framework for comparing classical reservoir computing (with a static linear readout), an attention-enhanced reservoir (with a dynamic readout), and transformer architectures on a character-level Shakespeare corpus. By systematically varying model sizes and measuring cross-entropy, N-gram Overlap, and time costs, we revealed how each architecture scales in performance and resource usage. Transformers excel overall, solidifying their state-of-the-art machine learning status, though they are more demanding in terms of computation time due to the quadratic complexity of the sequence length. Traditional reservoirs, while limited in their ability to capture complex dependencies, show surprising good results for their simple buildup. Attention-enhanced reservoirs strike a balance between efficiency and performance, outperforming traditional reservoirs and even getting close to the performance of Transformer models while having a negotiable complexity depending on the sequence length.
Acknowledgements.
This study was supported in part by JSPS KAKENHI (JP22H05195, JP25H01129) and JST CREST (JPMJCR24R2) in Japan.